OpenStudy (anonymous):

how to calculate 41% of 28 feet,please show steps.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Rewrite %41 as \[\frac{ 41 }{ 100 }\] Multiply it with 28 After solving that, we get 11.48 as result

OpenStudy (anonymous):

\[\frac{ 41 * 28 }{ 100 }\]

OpenStudy (anonymous):

yes the aswer is 41*28/100

OpenStudy (anonymous):

got it, thanks

OpenStudy (anonymous):

I just drop the percent sign and make it a decimal, then move it over 2 places to the left. Ex: .41*28=11.48